OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 2x + sin x

OpenStudy (anonymous):

it looks like a different problem

OpenStudy (anonymous):

cos(2x)=cos²x-sin²x

OpenStudy (anonymous):

and then sin x =

OpenStudy (anonymous):

cos^2x-sin^2x=1 ... therefore, 1+sinx=0 ... sinx=-1

OpenStudy (anonymous):

@Eng.Mido \[ \cos^2x +\sin^2x = 1 \]Not minus...

OpenStudy (anonymous):

so the answer woud be 1- 2 sin^2 x + sin x

OpenStudy (anonymous):

thanks ... I felt that sth was wrong too. :D

OpenStudy (anonymous):

cos2x=cos^2 x -sin^2x sin2x=2sinxcosx

OpenStudy (anonymous):

\[ \begin{split} \cos (2x) + \sin (x) &= \cos^2(x) - \sin^2(x) + \sin(x) \\ &= (1-\sin^2(x))-\sin^2(x) + \sin(x) \\ &= 1 - 2\sin^2(x) + \sin(x) \end{split} \]

OpenStudy (anonymous):

u could then take sinx as a common factor and solve the quadratic equation!!!

OpenStudy (anonymous):

Hmmm, actually we don't really need to use Pythagorean theorem here.

OpenStudy (anonymous):

@Eng.Mido Seems obvious to me that the roots are \(\pi / 2\) etc.

OpenStudy (anonymous):

Hopefully \[ 1 - 2\sin(x) \sin(x)+\sin(x) \]Is good enough?

OpenStudy (anonymous):

yep i think thats it

OpenStudy (anonymous):

@wio My reply was very clear ... As I used the word "Could"!!